pH of a solution at 250celsius is 3 if the pH  is to be doubled , then hydronium ion concentration should be

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
)]-7)and we know that at 25 degree C, pH+pOH=14,so calculate pH=14-pOH=14-[-log(2*10-7Mso pOH=-log(2*10-7 MAs per the questuion, OH-=2*10-7The hydroxide ions for pure water at 25C=10 RegardsArun(askIITians forum expert)